OpenStudy (anonymous):

The value of a car over several years is shown below: Age (years) Value (dollars) 0 $17,000 1 $15,000 2 $13,250 3 $11,700 4 $10,400 Find a model equation using a graphing technology tool to represent the age at any time.

OpenStudy (anonymous):

y =-6.02x + 10 y =-1650x + 16770 y =-1650x + 18420 y =-1650x + 17000